# Quantifying the Gain in Weak-to-Strong Generalization

### v1

@@ 1 Introduction
Towards answering this question, we show (Theorem 1) that the true error of a strong model trained on weak labels is smaller than the error of the weak model, by at least the error of the strong model on the weak labels itself. We call this latter quantity the misfit between the weak and strong model. Our result can be stated as the following simple principle:
@@
Intuitively, the misfit quantifies the erroneous knowledge that the strong model does not obtain from the weak model, and hence also the amount that the strong model improves over the weak model.
@@ 2.1 Related Work
The weak-to-strong generalization paradigm is perhaps most closely related to the teacher-student model of training [LA16, TV17] (sometimes also referred to as knowledge distillation [HVD15, GYMT21]), where a student model (typically smaller) is trained using data labeled by a teacher model (typically larger), and possibly some additional ground-truth data. The remarkable phenomenon of the student model outperforming the teacher has been observed in many works [BCNM06, HVD15, FLT+18]. Most relevant to us are formulations where the student model is equally [FLT+18] or more powerful [XLHL20] than the teacher model. There has been theoretical work explaining superior generalization in the setting where the student and teacher models are equally powerful, e.g., the work of [MFB20] where the student also has access to ground-truth labels, or the work of [WSCM20], where the model is trained on its own labels (known as self-training) with a regularization loss. In contrast, our work aims to explain why we might expect such behavior when the student is more powerful, and also doesn’t have any access to ground-truth labels.
@@ 4 Main Proof Technique
$\lambda(f\circ h_{s})(x)+(1-\lambda)(g\circ h_{s})(x)$ $=\lambda f(h_{s}(x))+(1-\lambda)g(h_{s}(x))$
$=p(h_{s}(x))=(p\circ h_{s})(x),$
@@ 5 Experiments
We perform experiments on synthetically generated data as well as real-world molecular prediction datasets to verify the guarantees on weak-to-strong generalization given by our theorems.
@@ 5.1 Synthetic Experiments
We set the target data representation $h^{\star}:\mathbb{R}^{8}\to\mathbb{R}^{16}$ to be a randomly initialized 5-layer multi-layer perceptron (MLP) with ReLU activations, with input dimension 8 and hidden layer dimension 16. The class $\mathcal{F}_{s}$ of finetuning tasks from which the strong model (as well as the weak model) learns is simply the class of linear functions from $\mathbb{R}^{16}\to\mathbb{R}$; $\mathcal{F}_{s}$ is thus a convex set (see Appendix C for instances where $\mathcal{F}_{s}$ is a non-convex set). The marginal data distribution $\mathcal{P}$ in our experiments is always $\mathcal{N}(0,\sigma^{2}I)$. To ensure that the data is well-spread, we set $\sigma=500$.
@@ 6 Conclusion
Our work has natural limitations. Our theorems notably do not apply when the set from which the strong model learns the finetuning task is not convex. Nevertheless, our experiments in Appendix C do suggest that our results should (at least qualitatively) hold even beyond the convex case. Our work also does not address classification tasks, and it would be interesting to see if similar results could be obtained for more general loss functions. Finally, while we do demonstrate results on real-world datasets, we anticipate that significantly larger-scale experiments on regression datasets used to train modern AI models will yield further interesting insights.
@@
This work is supported by Moses Charikar and Gregory Valiant’s Simons Investigator Awards.
@@ Appendix B Heuristic to Choose Among Weak Models
Appendix C Non-convex $\mathcal{F}_{s}$
@@
Appendix D Implementation details

### v2

@@ 1 Introduction
Towards answering this question, we show (Theorem 1) that in the concrete setting of regression, the true error of a strong model trained on weak labels is smaller than the error of the weak model, by at least the error of the strong model on the weak labels itself. We call this latter quantity the misfit between the weak and strong model. Our result can be stated as the following simple principle:
@@
Intuitively, the misfit quantifies the erroneous knowledge that the strong model does not obtain from the weak model, and hence also the amount that the strong model improves over the weak model. We note that the work of [BIK+23] does empirically show that the performance gain of the strong model scales directly with its misfit (or disagreement) with the weak model; our result thereby provides a precise quantification of this observation.
@@ 2.1 Related Work
The weak-to-strong generalization paradigm is perhaps most closely related to the teacher-student model of training [LA16, TV17] (sometimes also referred to as knowledge distillation [HVD15, GYMT21]), where a student model (typically smaller) is trained using data labeled by a teacher model (typically larger), and possibly some additional ground-truth data. The remarkable phenomenon of the student model outperforming the teacher has been observed in many works [BCNM06, HVD15, FLT+18]. Most relevant to us are formulations where the student model is equally [FLT+18] or more powerful [XLHL20] than the teacher model. There has been theoretical work explaining superior generalization of the student in some specialized settings, e.g., the work of [MFB20] where the student also has access to ground-truth labels, or the work of [WSCM20], which operates under a certain expansion criterion and consistency loss. In contrast, our work does not assume that the student model has any access to ground-truth labels, and also does not incorporate a regularization term in the objective. Finally, the conceptual insight in our work about the performance gain of the student model scaling with its disagreement with the teacher model is closely related to the theory of generalization bounds based on disagreement between different classifiers [DLM01, WZ17, YHY+19].
@@ 4 Main Proof Technique
$\lambda(f\circ h_{s})(x)+(1-\lambda)(g\circ h_{s})(x)$ $=\lambda f(h_{s}(x))+(1-\lambda)g(h_{s}(x))=p(h_{s}(x))=(p\circ h_{s})(x),$
@@ 5 Experiments
We perform experiments [^4] https://github.com/chogba/wtsg-regression on synthetically generated data as well as real-world molecular prediction and natural language datasets to verify the guarantees on weak-to-strong generalization given by our theorems. The results for the natural language tasks are given in Appendix C.
@@ 5.1 Synthetic Experiments
We set the target data representation $h^{\star}:\mathbb{R}^{8}\to\mathbb{R}^{16}$ to be a randomly initialized 5-layer multi-layer perceptron (MLP) with ReLU activations, with input dimension 8 and hidden layer dimension 16. The class $\mathcal{F}_{s}$ of finetuning tasks from which the strong model (as well as the weak model) learns is simply the class of linear functions from $\mathbb{R}^{16}\to\mathbb{R}$; $\mathcal{F}_{s}$ is thus a convex set (see Appendix D for instances where $\mathcal{F}_{s}$ is a non-convex set). The marginal data distribution $\mathcal{P}$ in our experiments is always $\mathcal{N}(0,\sigma^{2}I)$. To ensure that the data is well-spread, we set $\sigma=500$.
@@ 6 Conclusion
Our work has natural limitations. Our theorems notably do not apply when the set from which the strong model learns the finetuning task is not convex. Nevertheless, our experiments in Appendix D do suggest that our results should (at least qualitatively) hold even beyond the convex case. Our work also does not address classification tasks (see also Appendix E), and it would be interesting to see if similar results could be obtained for more general loss functions. Finally, while we do demonstrate results on real-world datasets, we anticipate that significantly larger-scale experiments on regression datasets used to train modern AI models will yield further interesting insights.
@@
This work is supported by Moses Charikar and Gregory Valiant’s Simons Investigator Awards. The authors would like to thank Percy Liang and Tengyu Ma for helpful discussions, and also the anonymous reviewers for engaging with the paper and pointing out references.
@@ Appendix B Heuristic to Choose Among Weak Models
Appendix C Experiments on NLP Datasets
We include additional experimental results on NLP regression tasks corresponding to two publicly available natural language datasets. The gain-misfit plots in these experiments largely follow the same trend as in Section 5.1 and Section 5.2.
C.1 Essay Scoring
We use the dataset from the Kaggle competition “Feedback Prize - English Language Learning" [FMB+22] conducted in 2022. Here, the task is to assign a score to essays written by 8th-12th grade English Language Learners (ELLs). An essay is assigned a score from 1-5 on each of 6 rubrics: cohesion, syntax, vocabulary, phraseology, grammar, conventions. Thus, a separate regression task can be obtained by treating each of these scores as the target label. The data source provides labels only for the train split; hence, we randomly do a 70-30 split of this data into train and test data.
We fix the strong model representation $h_{s}$ to be a standard pretrained bert-base-uncased model comprising of 110M parameters. We vary the weak model representation $h_{w}$ amongst different pretrained miniature BERT architectures [TCLT19]. Namely, we vary the hidden size $H\in\{128,256,512,768\}$ and the number of transformer layers $L\in\{2,4,6,8\}$ (the number of attention heads is always fixed at $H/64$). Pretrained weights for each of these is available online.
We perform the entire weak-to-strong supervision experiment identically as in Section 5.2—the gain-misfit plots are shown in Figure 5.
(a) cohesion
(b) syntax
(c) vocabulary
(d) phraseology
(e) grammar
(f) conventions
Figure 5: Results on the Essay Scoring dataset. Each plot corresponds to the task of predicting the score based on a different rubric.
C.2 French Public Service Reviews
We also perform experiments on a dataset of public service reviews in French language collected from Google Maps and Trustpilot which accompanies the blog post of [R&D22] available at https://lajavaness.medium.com/regression-with-text-input-using-bert-and-transformers-71c155034b13. The task is to assign a score in 0-4 to the reviews, indicating whether it is negative or positive. The data source provides train and test splits.
We use pretrained models from the CamemBERT [MMS+20] and FlauBERT [LVF+20] family for this experiment. We fix $h_{s}$ to be camembert-large (335M parameters), and range $h_{w}$ amongst camembert-base, camembert-base-ccnet, camembert-base-wikipedia-4gb, camembert-base-oscar-4gb, camembert-base-ccnet-4gb, flaubert-small-cased, flaubert-base-cased and flaubert-base-uncased. The results can be seen in Figure 6.
Figure 6: Results on the French Reviews dataset.
Appendix D Non-convex $\mathcal{F}_{s}$
@@
Appendix E Discussion Regarding Extensions to Classification
While our weak-to-strong generalization theory is tailored to regression, we do believe that the qualitative conclusion of our work should also extend to other settings, including classification. In particular, as evidenced by numerous plots on student-supervisor disagreement in [BIK+23], the performance gain in classification ought also be (mathematically) characterized, at least to some extent, by the misfit between the weak and strong model; however, additional regularization terms in the objective (along with other error terms) seem to be necessary. This is supported by Figure 8 in [BIK+23], which shows that adding auxiliary losses in the objective helps reduce student-supervisor agreement as compared to naively finetuning (without any auxiliary loss) to the weak labels, and thereby improves the perfomance of the strong student model.
Of particular relevance here is also the study in Appendices E.1, E.2 in [BIK+23], which studies the qualitative structure of weak-to-strong disagreement in the setting of classification. Namely, the question studied there is: do different forms of disagreement between the strong and weak model (e.g., random errors vs correlated errors), for the same weak model accuracy, lead to differing weak-to-strong generalization? While this nuance arises in the setting of classification, it may be worth noting that in the setting of regression that we consider, projection onto the convex set only results in movement towards the true function, assuming realizability—in this sense, any misfit is only ever “in the correct direction”, and its full quantitative benefit is realized. Thus, such a qualitative difference among different types of misfits does not manifest for us. Nevertheless, while it is true that in other settings, the nature of disagreement might matter, the general principle of decreasing student-supervisor imitation (alternatively, increasing misfit) to foster weak-to-strong generalization, either via auxiliary losses or otherwise, does constitute a significant theme in the work of [BIK+23].
Appendix F Implementation Details